Morton Company’s contribution format income statement for last month is given below:

Sales (44,000 units × $23 per unit)

$

1,012,000

Variable expenses

708,400

Contribution margin

303,600

Fixed expenses

242,880

Net operating income

$

60,720

The industry in which Morton Company operates is quite sensitive to cyclical movements in the economy. Thus, profits vary considerably from year to year according to general economic conditions. The company has a large amount of unused capacity and is studying ways of improving profits.

Required:

1. New equipment has come onto the market that would allow Morton Company to automate a portion of its operations. Variable expenses would be reduced by $6.90 per unit. However, fixed expenses would increase to a total of $546,480 each month. Prepare two contribution format income statements, one showing present operations and one showing how operations would appear if the new equipment is purchased.

2. Refer to the income statements in (1). For the present operations and the proposed new operations, compute (a) the degree of operating leverage, (b) the break-even point in dollar sales, and (c) the margin of safety in dollars and the margin of safety percentage.

3. Refer again to the data in (1). As a manager, what factor would be paramount in your mind in deciding whether to purchase the new equipment? (Assume that enough funds are available to make the purchase.)

4. Refer to the original data. Rather than purchase new equipment, the marketing manager argues that the company’s marketing strategy should be changed. Rather than pay sales commissions, which are currently included in variable expenses, the company would pay salespersons fixed salaries and would invest heavily in advertising. The marketing manager claims this new approach would increase unit sales by 30% without any change in selling price; the company’s new monthly fixed expenses would be $387,596; and its net operating income would increase by 20%. Compute the company's break-even point in dollar sales under the new marketing strategy.
Answer #1

1) Before purchasing the equipment

Particulars Amount
Sales 1012000
Less: Variable expenses (708400)
Contribution 303600
Less: Fixed cost (242880)
Profit 60720

If the equipment is purchased,

Particulars Amount
Sales 1012000

Less: Variable expenses

{ 708400/44000=16.1

16.1 - 6.9 = 9.2

44000*9.2 = 404800 }

(404800)
Contribution 607200
Less : Fixed cost (546480)
Profit 60720

2)

Particulars Present operations Proposed new operations

a) Degree of operating leverage

(Contribution/EBIT)

5

(303600/60720)

10

(607200/60720)

b) Break even point in sales

(Fixed cost/P V ratio)

P V ratio = contribution/sales *100

809600

(242880/0.3)

( P V ratio = 303600/1012000*100 = 30% )

910800

(546480/0.6)

( P V ratio = 607200/1012000*100 = 60% )

c) Margin of safety

(Total sales - Break even sales)

202400

(1012000-809600)

101200

(1012000-910800)

Margin of safety percentage

( Margin of safety/sales*100)

20%

(202400/1012000*100)

10%

(101200/1012000*100)

3) The factor that would be paramount in our mind in deciding whether to purchase the new equipment is Cyclical movements in the economy itself. It is an external factor. It cannot be controlled easily. All the other factors are internal to the business.

4)As per the new marketing strategy,

Sales in units = 44000*130/100

= 57200 units

Therefore,

Sales = 57200 * 23

= $1315600

Profit

(60720*120/100)

72864
Add : Fixed cost 387596
Contribution 460460

P V ratio = Contribution/Sales*100

= 460460/1315600*100

= 35%

Break even point in sales = Fixed cost/P V ratio

= 387596/0.35

=$1107417